OpenStudy (anonymous):

x² + 5x - 36 Which binomial is a factor of the polynomial?

jimthompson5910 (jim_thompson5910):

Hint: find two numbers that multiply to -36 and add to 5

OpenStudy (anonymous):

i cant figure it out

OpenStudy (anonymous):

i thought it was (x+6) but thats not it

jimthompson5910 (jim_thompson5910):

list out the possible ways to factor -36

jimthompson5910 (jim_thompson5910):

add up each factor pair and see which result gives you 5

OpenStudy (anonymous):

and how would i find the factor pair

jimthompson5910 (jim_thompson5910):

for instance, one way to factor -36 is to write it as -2*18 notice how -2 + 18 = 16 so the two numbers aren't -2 and 18 so you have to keep looking

jimthompson5910 (jim_thompson5910):

36 = -3*12 -3 + 12 = 9 so that doesn't work either

jimthompson5910 (jim_thompson5910):

you keep going until you have a pair that add to 5 or until you've checked every pair

OpenStudy (anonymous):

only other one i can think of is 6*6

jimthompson5910 (jim_thompson5910):

6 + 6 = 12, so that doesn't work -6 + 6 = 0, that doesn't work

jimthompson5910 (jim_thompson5910):

what else haven't you tried

OpenStudy (anonymous):

4*9

jimthompson5910 (jim_thompson5910):

which one is negative

OpenStudy (anonymous):

-4

jimthompson5910 (jim_thompson5910):

-4*9 = -36 -4 + 9 = 5 both work

jimthompson5910 (jim_thompson5910):

this means x^2 + 5x - 36 factors to (x - 4)(x + 9)

OpenStudy (anonymous):

ohhh ok thank you so much

jimthompson5910 (jim_thompson5910):

yw